OpenStudy (anonymous):

How do the graphs of y = 1/x and y=5/x+6 compare?

OpenStudy (anonymous):

Now we can compare @SmokeysTheName

OpenStudy (smokeysthename):

idek how to do this

OpenStudy (anonymous):

Thats number 14 on the workpad. And yeah me neither thats why i posted it

OpenStudy (smokeysthename):

alright

OpenStudy (anonymous):

Do you understand 15?

OpenStudy (smokeysthename):

not at all.. lemme post that question.

OpenStudy (anonymous):

Okay great

OpenStudy (smokeysthename):

@phi can you help us with this?

OpenStudy (anonymous):

@master43 can you help us?

OpenStudy (anonymous):

^~^

OpenStudy (anonymous):

Yes , No , maybe so ?

OpenStudy (smokeysthename):

please?

OpenStudy (anonymous):

Can you help us @phi

OpenStudy (smokeysthename):

@phi i think its like a vertical shift to the left 6 and a horizontal shrink by 5.. is that right?

OpenStudy (phi):

First look at y = 1/x if x is 0 then y= 1/0 which is undefined. But think like this: instead of 0, let x be a tiny number like 1/100000 when we divide 1/ (1/100000) we get 100000 (a big number) the point is as you get close to x=0 y gets really big when x=1, y =1/1 = 1 when x is really big y = 1/10000 which is close to 0 so here are 3 points (0, infinity), (1,1), and (infinity, 0) it looks like this